# Introducing Delays in Multi-Agent Path Finding

### v3

@@ 1 Introduction
Multi-Agent Path Finding (MAPF) is a problem in Artificial Intelligence (AI) that asks to find non-colliding paths for a group of agents moving on a graph (Stern et al. 2019; Salzman and Stern 2020). Applications vary from autonomous warehouse management (Wurman, D’Andrea, and Mountz 2008) and factory pipe routing (Belov et al. 2020) to rail planning (Li et al. 2021) and swarm robotics (Ramaithitima et al. 2016; Araki et al. 2017). Although MAPF is known to be generally an intractable problem (Yu 2016; Banfi, Basilico, and Amigoni 2017; Nebel 2020; Geft 2023), recent algorithms can scale to thousands of agents, e.g., (Li et al. 2022; Li, Ruml, and Koenig 2021; Okumura 2023). A limiting aspect of these algorithms is the simplifying assumption that, at deployment, agents can synchronously execute a plan. In reality, however, it is common for agents to fall out of sync, e.g., due to delays or model uncertainty. Such incidents may cause the plan to no longer be valid (non-colliding), in which case we must either compute a new plan or repair the old plan quickly. This is challenging since replanning faces the same difficulties as the original MAPF problem, and plan repair is shown to be as difficult as plan generation itself (Nebel and Koehler 1995).
@@ 3 Computational Complexity of ACID
We are now ready to establish the complexity of ACID. The complete proof can be found in the technical appendix. For background on hardness of approximation see (Ausiello et al. 1999).
@@ 4.2 Improved Constrained Graph
As we prove in the technical appendix, using ICG instead of CG is sound and complete, in the following sense.
@@ 5 Experimental Evaluation
Figure 6: Global averages over all tested instances. CG and ICG lines are coincident in Figs. 5(a)-5(d), and 5(g).
@@ 5.2 Computation Time
10 Success Rate (%) 1 32 32 99 98 98 32 95 100
Computation Times (s) 5.3 5.3 1.6 2.1 1.7 5.4 3.7 4.5
Added SOC 6.8 6.8 586.7 (167.7) 346.3 (80) 345.5 (78.5) 230.4 (65.6) 136.1 (98.7) 140.2 (102.4)
50 Success Rate (%) 0 0 0 100 100 100 31 100 100
Computation Times (s) 1.8 2.8 2.1 7.1 5.2 5.5
Added SOC 659.5 (318.3) 758 (295.9) 757.4 (313.7) 336.3 (138.1) 555.8 (370.2) 566.8 (378.9)
Table 4: Experimental results for $10$ and $50$ simultaneous delays. Anytime algorithms (Anytime-EECBS and MAPF-LNS2) report two values for Added SOC: the initial solution and the best solution (in parentheses). Computation times are provided for the initial solutions.
The average computation times of every MAPF solver over all of the instances are shown in Figs. 5(b)5(e), and 5(h). The three lines correspond to OG (blue), CG (orange), and ICG (green). The most obvious computation time improvements are seen in Fig. 5(b) which show that not only can CBS on CG and ICG scale to $1000$ agents, but it also produced very low computation times compared to OG. Similarly, MAPF-LNS2 on CG and ICG performed faster than on OG (see Fig. 5(h)). Anytime-EECBS on ICG improved its computation times over both CG and OG, especially for large number of agents (see Fig. 5(e)).
@@ 5.3 Success Rate
The average success rates of every MAPF solver over all the instances are shown in Figs. 5(a)5(d), and 5(g). The results show that CG and ICG perform better than OG across all values of $n$ for both CBS and MAPF-LNS2 while performing equally as well for Anytime-EECBS. CBS improved the most, which succeeded over $80\%$ of the time for all $n$ with CG and ICG but only scaled to $400$ agents on OG.
@@ 5.4 Added SOC
The average added SOC of every MAPF solver over all the maps are shown in Figs. 5(c)5(f), and 5(i). The results show that both Anytime-EECBS and MAPF-LNS2 generally provided shorter solutions when using CG and ICG compared to OG (see Figs. 5(f) and 5(i)). Replanning with CBS on CG and ICG generally results in very small SOC additions (about $4\ll n$ as per Remark 3) compared to Anytime-EECBS and MAPF-LNS2 (see Fig. 5(c)).
@@ 5.5 Experimental setup – multiple delays
Our setup for introducing multiple delays is identical to that of Section 5.1 with some key exceptions: we introduced multiple simultaneous delays, we tested on $3$ maps (Figs. 4(e)4(f), and 4(h)), we only considered instances with $600$ agents, our timeout was one minute, and we considered the anytime properties of Anytime-EECBS and MAPF-LNS2. Table 4 report our results for $10$ and $50$ delay introductions.

### v4

@@ 1 Introduction
Multi-Agent Path Finding (MAPF) is a problem in Artificial Intelligence (AI) that asks to find non-colliding paths for a group of agents moving on a graph (Stern et al. 2019; Salzman and Stern 2020). Applications vary from autonomous warehouse management (Wurman, D’Andrea, and Mountz 2008) and factory pipe routing (Belov et al. 2020) to rail planning (Li et al. 2021) and swarm robotics (Ramaithitima et al. 2016). Although MAPF is known to be generally an intractable problem (Yu 2016; Banfi, Basilico, and Amigoni 2017; Nebel 2020; Geft 2023), recent algorithms can scale to thousands of agents, e.g., (Li et al. 2022; Li, Ruml, and Koenig 2021; Okumura 2023). A limiting aspect of these algorithms is the simplifying assumption that, at deployment, agents can synchronously execute a plan. In reality, however, it is common for agents to fall out of sync, e.g., due to delays or model uncertainty. Such incidents may cause the plan to no longer be valid (non-colliding), in which case we must either compute a new plan or repair the old plan quickly. This is challenging since replanning faces the same difficulties as the original MAPF problem, and plan repair is shown to be as difficult as plan generation itself (Nebel and Koehler 1995).
@@ 3 Computational Complexity of ACID
We are now ready to establish the complexity of ACID. The complete proof can be found in Section 7. For background on hardness of approximation see (Ausiello et al. 1999).
@@ 4.2 Improved Constrained Graph
As we prove in Section 7, using ICG instead of CG is sound and complete, in the following sense.
@@ 5 Experimental Evaluation
Figure 6: Global averages over all tested instances. CG and ICG lines are coincident in Figs. 6(a)-6(d), and 6(g).
@@ 5.2 Computation Time
10 Succ. Rate (%) 1 32 32 99 98 98 32 95 100
Comp. Times (s) 5.3 5.3 1.6 2.1 1.7 5.4 3.7 4.5
Added SOC 7 7 587 (168) 346 (80) 346 (79) 230 (66) 136 (99) 140 (102)
50 Succ. Rate (%) 0 0 0 100 100 100 31 100 100
Comp. Times (s) 1.8 2.8 2.1 7.1 5.2 5.5
Added SOC 660 (318) 758 (296) 757 (314) 336 (138) 556 (370) 567 (379)
Table 4: Experimental results for $10$ and $50$ simultaneous delays. Anytime algorithms report two values for Added SOC: the initial solution and the best solution (in parentheses). Computation times are provided for the initial solutions.
The average computation times of every MAPF solver over all of the instances are shown in Figs. 6(b)6(e), and 6(h). The three lines correspond to OG (blue), CG (orange), and ICG (green). The most obvious computation time improvements are seen in Fig. 6(b) which show that not only can CBS on CG and ICG scale to $1000$ agents, but it also produced very low computation times compared to OG. Similarly, MAPF-LNS2 on CG and ICG performed faster than on OG (see Fig. 6(h)). Anytime-EECBS on ICG improved its computation times over both CG and OG, especially for large number of agents (see Fig. 6(e)).
@@ 5.3 Success Rate
The average success rates of every MAPF solver over all the instances are shown in Figs. 6(a)6(d), and 6(g). The results show that CG and ICG perform better than OG across all values of $n$ for both CBS and MAPF-LNS2 while performing equally as well for Anytime-EECBS. CBS improved the most, which succeeded over $80\%$ of the time for all $n$ with CG and ICG but only scaled to $400$ agents on OG.
@@ 5.4 Added SOC
The average added SOC of every MAPF solver over all the maps are shown in Figs. 6(c)6(f), and 6(i). The results show that both Anytime-EECBS and MAPF-LNS2 generally provided shorter solutions when using CG and ICG compared to OG (see Figs. 6(f) and 6(i)). Replanning with CBS on CG and ICG generally results in very small SOC additions (about $4\ll n$ as per Remark 3) compared to Anytime-EECBS and MAPF-LNS2 (see Fig. 6(c)).
@@ 5.5 Experimental Setup – Multiple Delays
Our setup for introducing multiple delays is identical to that of Section 5.1 with some key exceptions: we introduced multiple simultaneous delays, we tested on $3$ maps (Figs. 5(e)5(f), and 5(h)), we only considered instances with $600$ agents, our timeout was one minute, and we considered the anytime properties of Anytime-EECBS and MAPF-LNS2. Table 4 report our results for $10$ and $50$ delay introductions.